OpenStudy (anonymous):

Suppose you have 54 feet of fencing to enclose a rectangular dog pen. The function a=27x-x^2, where x=width, gives you the area of the dog pen in square feet. What width gives you the maximum area? Round to the nearest tenth as necessary.

OpenStudy (amistre64):

what is the axis of symmetry for your quadratic?

OpenStudy (amistre64):

or if you know the calculus, define the derivative to be zero

OpenStudy (anonymous):

This is for algebra, so to be honest, I'm not sure. My friend was looking at this problem, I was intrigued by it, so I tried it out.

OpenStudy (amistre64):

do you know what the shape of the graph would look like?

OpenStudy (amistre64):

there is also a factoring way to approach it; define the midpoint between the zeros

OpenStudy (anonymous):

Well I do know that the answer is supposed to be width= 13.5 ft; area=182.3 ft squared

OpenStudy (anonymous):

I'm real sorry, I didn't really know much background before asking this question, it's alright if you don't want to solve it

OpenStudy (amistre64):

since this is algebra; lets define the values of x for which it is zero 27x-x^2 = 0 x (27-x) = 0 now, we should know that when we multiply anything by 0 we get 0; so this tells us that either: x = 0, 0 (27-0) = 0 or x = 27, 27 (27-27) = 0

OpenStudy (amistre64):

a quadratic will have its minimum or maximum value (depending on how it opens) at the middle point between the zeros halfway between 0 and 27 is: 13.5

OpenStudy (anonymous):

Hmmm.... so assuming that similar problems were taught in a school lesson, they would first have to instruct the class that the minimum/maximum is at the middle point between the zeros

OpenStudy (amistre64):

in algebra class, they go over the properties of a quadratic equation; so this would have been reviewed at some point during the class yes

OpenStudy (amistre64):

some quadratics never touch the x axis, so there are a few ways to approach this; this was just an easy one to assess by facoting

OpenStudy (amistre64):

*factoring

OpenStudy (anonymous):

So then what steps would I use to go about figuring out the maximum area? THANK YOU SO MUCH BY THE WAY

OpenStudy (amistre64):

well, once you know the value for x, plug it in and work the 4th grade math skills ...

OpenStudy (anonymous):

Alright, much appreciated, I really have to start easing back into math before school start ;)

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks!